Journalist Report – April 27th

Crew #: 335
Position: Crew Journalist
Report prepared by: Matteo Calore
Date: 27APR2026
Sol: Day 8

Report Title: Windy Focus

Today was a long day. We woke up, as usual, between six and seven, and had a quick breakfast before moving into our regular morning meeting, pleased to see that we are managing our resources well. The challenge of living in a space with limited resources is always interesting, but for some of us it is not new, so it comes quite naturally.

As the days go by, each of us is increasingly focused on completing our own projects, so interactions have become fewer. From the initial desire to get to know one another, we now feel the pressure of time and the need to bring to completion what we started. Overall, we are all making good progress, and aside from a few interruptions, we remain well organized.

The desert and the isolation help us stay more focused, even though some of us are finding it more difficult, especially due to trouble sleeping at night. Some are particularly affected by the dry desert air. Still, we are gradually adapting to this way of life, and being surrounded by the beauty of these hills—where red and white dominate, touched by the intense sunlight—helps us stay concentrated.

Sol Summary – April 27th

Sol Summary Report

Crew #: 335
Position: Crew Artist
Report prepared by: Lot Brugge
Date: 27APR2026
Sol: Day 8
Summary Title: Windy
Mission Status: Towards the end

Sol Activity Summary
We woke up around 07:00 and had a quick breakfast. One crew member continued to rest due to an accident the previous day and is not yet in good condition, so additional recovery time is needed to continue the final days of work.

The rest of the crew began the day with the usual morning meeting to review activities and overall status. Water consumption remains stable. Some crew members completed reports that had been delayed, as the previous day’s schedule was interrupted by a visiting crew of journalists from England, preventing us from fully carrying out planned simulation activities.

After the meeting, each crew member focused on their individual projects. Charlotte continued working on her clay project, while the film crew documented activities and organized interviews with each crew member.

In the afternoon, the film crew conducted an interview with the Commander and later performed a short EVA around the station to document MDRS at sunset. The rest of the crew remained inside: Charlotte continued her work, and the Commander rested in the living area.

Anomalies in Work
The planned EVA destination was changed and postponed to the following day due to strong wind conditions and the health status of one crew member.

Weather
The morning was sunny and mild. In the afternoon, wind increased significantly, and light rain occurred around sunset.

Crew Physical Status
One crew member continues to experience a persistent cough. Another is recovering from an accident and requires rest.

EVA
The filmmakers conducted a short EVA to capture footage of MDRS at sunset.

Comms Window Closed 08Apr2026

Mission Support is signing off.

Report status for sol 10:

  • Sol Summary – received
  • Operations Report – received
  • Greenhab Report – received
  • Journalist Report – received
  • Astronomy Report – received
  • EVA Report – received
  • EVA Request(s) – EVA 17 and 18 approved
  • Daily photos – received

Best regards,
________________________
Sergii Iakymov – Director Mars Desert Research Station

AIorK4zJwLwPIWpaKeu3MS1SRDkfIkROXwfkNaEgVlcUFUQOHMM_jnKth8pJaFRg3ou53q1RY40muac

GreenHab Report – February 22nd

Report title: GreenHab Report
Crew #: 330
Position: Greenhab officer
Report prepared by: Félix Marguet
Date: 22-02-2026
Sol: 7

Environmental control (fan & heater): Yes
Average temperatures (last 24h): 75.7°F
Maximum temperature (last 24h): 85.5°F
Minimum temperature (last 24h): 67.9°F
Average humidity (last 24h): 19,4%
Maximum humidity (last 24h): 27.9%
Minimum humidity (last 24h): 13.2%

Hours of supplemental light: 5h30
Daily water usage for crops: 2,5gal
Daily water usage for research and/or other purposes: 0 gal
Water in the Tank (160 gal useful capacity): 150 gallons remaining
Time(s) of watering for crops: 11AM, 5PM
Changes to crops: N/A
Narrative: N/A
Harvest: (include which crop and mass in grams) none
Support/supplies needed : none

Supplemental Operations Report 21Feb2026

[title Supplemental Operations Report – February 21st]

Date: 02/21/2026
Name of person filing report: Anderson Wilder
Reason for Report: Routine
Non-Nominal Systems: Hab outer shell. Percy e-brake

Power system:
1) 200 hr maintenance (including oil, oil filter, and air filter change) expected routine time
2) Current generator hours – 1434.5

Propane Readings:
Station Tank: 54%
Director Tank: 55%
Intern Tank: 60%
Generator Tank: 71%
Propane refilled on 1/20/26

Water:
Hab Static Tank – 420 gallons
GreenHab – 200 gallons
Outpost tank – 400 gallons
Septic Tanks emptied: No

Rovers:
Sojourner rover used: No
Hours: 222.5
Beginning Charge: 100 %
Ending Charge: _ %
Currently Charging: Yes

Percy e-brake not working, have replacement, need to replace

Cars:
Hab Car used and why, where: To Hanksville for supplies
Crew Car used and why, where: To Hanksville for supplies, Grand Junction for crew pick up.
General notes and comments: Crew car driver’s side front ball joint is bad

Summary of Internet: Nominal.
EVA suits and radios:
Suits: Nominal.
Comms: Nominal.
T-Echo EVA-link: Nominal.

Campus:Nominal.
Summary of Hab Operations: Nominal.
Summary of GreenHab Operations: Nominal.
Summary of SciDome Operations: Nominal.
Summary of Observatories Operations: Nominal.
Summary of RAM Operations: Cleaned. Nominal.
Summary of Outpost Operations: Nominal.
Summary of Health and Safety Issues: Nominal.

Supplemental Operations Report 15Feb2025

[title Supplemental Operations Report – February 15th]

Date: 02/15/2025
Name of person filing report: Anderson Wilder
Reason for Report: Routine
Non-Nominal Systems: Hab outer shell. Percy e-brake

Power system:
1) 200 hr maintenance (including oil, oil filter, and air filter change) expected routine time
2) Current generator hours – 1419.2

Propane Readings:
Station Tank: 61%
Director Tank: 64%
Intern Tank: 65%
Generator Tank: 77%
Propane refilled on 1/20/26

Water:
Hab Static Tank – 550 gallons
GreenHab – 220 gallons
Outpost tank – 500 gallons
Septic Tanks emptied: No

Rovers:
Sojourner rover used: Training
Hours: 222.5
Beginning Charge: 100 %
Ending Charge: _ %
Currently Charging: Yes

Percy e-brake not working, have replacement, need to replace

Cars:
Hab Car used and why, where: To Hanksville for supplies, Green River
Crew Car used and why, where: To Hanksville for supplies, Grand Junction for crew
General notes and comments: N/A

Summary of Internet: Nominal.
EVA suits and radios:
Suits: Nominal.
Comms: Nominal.
T-Echo EVA-link: Nominal.

Campus:Nominal.
Summary of Hab Operations: Nominal.
Summary of GreenHab Operations: Nominal.
Summary of SciDome Operations: Nominal.
Summary of Observatories Operations: Nominal.
Summary of RAM Operations: Cleaned. Nominal.
Summary of Outpost Operations: Nominal.
Summary of Health and Safety Issues: Nominal.

GreenHab Report – February 23rd

Report title: GreenHab Report
Crew #: 330
Position: Greenhab officer
Report prepared by: Félix Marguet
Date: 23-02-2026
Sol: 8

Environmental control (fan & heater): Yes
Average temperatures (last 24h): 75.5°F
Maximum temperature (last 24h): 86.12°F
Minimum temperature (last 24h): 68.6°F
Average humidity (last 24h): 22%
Maximum humidity (last 24h): 27.4%
Minimum humidity (last 24h): 14.3%

Hours of supplemental light: 5h30
Daily water usage for crops: 2,5gal
Daily water usage for research and/or other purposes: 0 gal
Water in the Tank (160 gal useful capacity): 146 gallons remaining
Time(s) of watering for crops: 11AM, 5PM
Changes to crops: I planted cucumber seeds in the big blue box.
Narrative: N/A
Harvest: (include which crop and mass in grams) none
Support/supplies needed : none

EVA Report – February 23rd

Report title: EVA Report

Crew #: 330

Position: Journalist

Report prepared by: Adrien Saugnac

Date: (23-02-2026)

Sol: 8

EVA #: 07

Purpose of EVA: Exploration and Photogrammetry experiment

The photogrammetry experiment aims to create 3D maps of the terrain around the MDRS. The quality and usefulness of the map will be assessed by other EVAs, using the map to orient themselves. This maps are created using Drone imagery, and this first EVA aims to map out a first point of interest outside of the station.

Start time: 9 : 55 AM

End time: 11 : 35 AM

Narrative: We took 18 minutes with Curiosity and Opportunity to go to Compass Rock because we had trouble finding the entry of Galileo Road 1104. Once we were on site, we started the drone and took pictures of the site, and we decided where the checkpoint of the orienteering race will be. The goal will be to evaluate the quality and usefulness of the 3D map.

At 11:10 AM, we headed back to the station and the EVA stopped at 11:35 AM without issues.

Destination per MDRS Map: Compass Rock

Destination Coordinates (use UTM WGS 84): 520000 E, 4252000 N

EVA Participants:

– EVA Leader : Adrien, Journalist
– Célia, Astronomer
– Lou, Engineer

Road(s) and routes per MDRS Map: Cow Dung Road towards North, Galileo Road 1104

Mode of travel: driving and walking

Operations Report – February 23rd

Report title: Operations Report

Crew #: 330

Position: Crew Engineer

Report prepared by: Lou Goiffon

Date: 23-02-2026

Sol: 8

Non-nominal systems: EVA Radio sets

Notes on non-nominal systems: Still some trouble with some radios or headsets, to be investigated by the crew

ROVERS

Spirit rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used : Yes

Hours (Before EVA): 256.1

Beginning charge: 100%

Ending charge: 43%

Currently charging: YES

Curiosity rover used: Yes

Hours (Before EVA): 347.2

Beginning charge: 100%

Ending charge: 44%

Currently charging: YES

Perseverance rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: YES

General notes on rovers:

Summary of Hab operations:

Water Use (see notes): 25.75 gal/d

Main tank level (remaining gallons, see notes): 197 gallons

Main water tank pipe heater (ON or OFF): ON

Main water tank heater (ON or OFF): ON

Toilet tank emptied (NO or YES): NO

Summary of internet: downlink 365.27 Mbits/s, uplink 16.44 Mbits/s, ping 25 ms

Summary of suits and radios:

– N°1 : Beginning charge 100%, ending charge 83%

– N°4 : Beginning charge 100%, ending charge 0% (the fan was still working a little bit at the end but it was not strong) it seems like there is a problem with battery, I will test it tomorrow

– N°6 : Beginning charge 100%, ending charge 86%

All the GPS beacons worked

Summary of GreenHab operations: (use information from the GreenHab report)

Water use: 2.5 gallons

Heater (ON or OFF): ON

Supplemental light (hours of operation): 6pm – 11:30pm

Harvest (name, weight in grams): 0

Summary of ScienceDome operations: Human factors experiments conducted

Dual split (Heat or AC, On or Off or Automatic): ON

Summary of RAM operations: Used the manual cart for the second EVA

Summary of any observatory issues: None

Summary of health and safety issues: None

Questions, concerns and requests to Mission Support: N/A